OpenStudy (anonymous):

Solve the following problem. It may be helpful to draw a chart on scrap paper to organize the information and write the equation. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. A clerk was asked to change a $10 bill. She returned 9 more dimes than nickels and twenty-one more quarters than dimes. How many coins of each did she return?

OpenStudy (texaschic101):

.10d + .05n + .25q = 10 d = n + 9 q = d + 21 or (n + 9) + 21, which is n + 30 so lets sub.. .10(n + 9) + .05n + .25(n + 30) = 10 .10n + .90 + .05n + .25n + 7.50 = 10 .40n + .90 + 7.50 = 10 .40n + 8.40 = 10 .40n = 10 - 8.40 .40n = 1.60 n = 1.60/.40 n = 4 d = n + 9 d = 4 + 9 d = 13 q = d + 21 q = 13 + 21 q = 34 check... .10d + .05n + .25q = 10 .10(13) + .05(4) + .25(34) = 10 1.30 + .20 + 8.50 = 10 10 = 10 (correct) 4 nickels, 13 dimes, and 34 quarters

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)